OpenStudy (anonymous):

Solve algebraically for the positive value of x, x ≠ 0, and check: 2x+5/7=1/x

OpenStudy (anonymous):

\[\frac{2x+5}{7}=\frac{1}{x}\] \[x(2x+5)=7\] \[2x^2+5x-7=0\] \[x=\frac{-5 \pm \sqrt{25-4(2)(-7)}}{2(2)}\] \[x=\frac{-5 \pm \sqrt{81}}{4}=\frac{-5 \pm 9}{4}=1, \frac{-7}{2}\] You only want the positive x, so discard x=-7/2

OpenStudy (anonymous):

oh i see. thank you

OpenStudy (anonymous):

(2x+5)/7 = 1/x (2x+5) x = 7 2x² + 5x - 7 = 0 ( x -1) ( 2x + 7) = 0 => x = 1 ( Omit x = -7/2)